How do you get the formula a+c=2b for the sides of a triangle in A.P?
 

Grade:12

1 Answers

Riddhish Bhalodia
8 years ago
OK
If the sides are in AP then they have a constant difference between consecutive sides and so we can write the side as
a ,b = a + d and c =a+2d where d is the constant difference
and hence
a+c =2b